OpenStudy (aonz):

Last question!!! log 1/5 (25)

OpenStudy (aonz):

check the identity by evaluating...

OpenStudy (aonz):

\[\huge \log_{\frac{ 1 }{ 5 }}25 \]

hartnn (hartnn):

use the same property i gave for last Q

Parth (parthkohli):

\(\frac{1}{5} = 5^{-1}\). What would you multiply to that to get \(5^2\)?

OpenStudy (aonz):



hartnn (hartnn):

yes, simplify numerator and denominator separately.

OpenStudy (agent0smith):

\[\large \log_{\frac{ 1 }{ 5 }} 5 = \frac{ \log_{5} 25 }{ \log_{5}\frac{ 1 }{ 5}} = \frac{ \log_{5} 5^2 }{ \log_{5} 5^{-1}}\]

Parth (parthkohli):

Why so serious?

Parth (parthkohli):

Just figure out what you multiply to \(5^{-1}\) to get \(5^2.\)

OpenStudy (aonz):

ahh @agent0smith explained it pretty good :)

OpenStudy (aonz):

@agent0smith How did you get the base to be 5? Like all of them?

OpenStudy (agent0smith):

Not sure what you mean... I just used base 5 because it makes it easy to solve.

Parth (parthkohli):

@AonZ Change of bases.

hartnn (hartnn):

you can choose any base for using that property,e,5,10,... thats why its unspecified in the property

OpenStudy (aonz):

so if you use change of base rule you can choose ANY base?

Parth (parthkohli):

Yes.

OpenStudy (aonz):

wow i reckon i understand logs a lot better now :P

OpenStudy (agent0smith):

^ correct @ParthKohli shouldn't it be "figure out what power you need to raise 5^-1 to, to get 25" as opposed to "Just figure out what you multiply to 5^−1 to get 5^2"?

Parth (parthkohli):

Yes, I am sorry.

Parth (parthkohli):

But you get the point.

OpenStudy (agent0smith):

Ah okay, cos i was wondering how this helps:\[5^{-1} \times 5^3 = 25\]compared to this \[(5^{-1})^{-2}=25\]

Parth (parthkohli):

:-)